# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

### v1

@@ Abstract.
We examine semiclassical measures for Laplace eigenfunctions on compact hyperbolic $(n+1)$-manifolds. We prove their support must contain the cosphere bundle of a compact immersed totally geodesic submanifold. Our proof adapts the argument of [DJ18] to higher dimensions and classifies the closures of horocyclic orbits using Ratner theory. An important step in the proof is a generalization of the higher-dimensional fractal uncertainty principle of Cohen [Co24] to Fourier integral operators, which may be of independent interest.
@@ 1.1. Main Results
Let $U_{1}^{-}$ be the vector field on $F^{*}M$ defined in (2.6). This vector field is the generator of the horocyclic flow $e^{sU_{1}^{-}}$.
It is well-known (see for example  [Zw12]*§§5.1-2) that if $\mu$ is a semiclassical measure, $\mu$ is invariant under the geodesic flow $\varphi_{t}:S^{*}M\rightarrow S^{*}M$ and $\operatorname{supp}\mu\subset S^{*}M$.
@@
We remark here that the presence of cosphere bundles of totally geodesic submanifolds in the higher-dimensional statement of both Theorem 1.2 and the main theorem of [ADM24] is intimately related to the failure of $1$-parameter unipotent flows on $F^{*}M$ to be uniquely ergodic. This should be juxtaposed with the $2$-dimensional case [DJ18], where the unique ergodicity of the horocycle flow is a key tool in concluding that semiclassical measures have full support.
@@ 1.2. Previous results
Developed in the 1970s and 80s, the Quantum Ergodicity theorem of Shnirelman [Sh74], Zelditch [Ze87], and Colin de Verdière [CdV89] implies that on negatively curved compact manifolds, a density-one sequence of Laplace eigenfunctions converges semiclassically to the Liouville measure.
@@
Theorems 1.1 and 1.2 exclude different semiclassical measures than the entropy bounds. This distinction is perhaps easiest to state in dimension $3$, where the existence of infinitely many commensurability classes of compact hyperbolic $3$-manifolds with no immersed, totally geodesic surfaces is known [MRBook, Theorem 9.5.1]. For such examples, Theorem 1.2 asserts that any semiclassical measure has full support, in contrast to the abundance of measures with entropies greater than $1$ with proper support. In higher dimensions, one can make similar statements, though a limiting factor is our understanding of the behavior of geodesic submanifolds in non-arithmetic manifolds (see §3.6). However, the entropy bounds exclude some measures that our results do not. One example is a semiclassical measure of the form $\mu=\alpha\mu_{L}+(1-\alpha)\mu_{0}$, where $\mu_{L}$ is the Liouville measure, $\mu_{0}$ is a $\delta$-measure on a closed geodesic, and $\alpha<1/2$. Such a measure clearly has full support, but has entropy $\alpha n<n/2$, and therefore is ruled out by [AN07].
Our work, specifically Theorem 1.1, builds off of [DJ18], wherein Dyatlov and Jin proved that all semiclassical measures on compact hyperbolic surfaces $M=\Gamma\backslash\mathbb{H}^{2}$ have full support. This result was previously generalized to compact surfaces with negative curvature by Dyatlov, Jin, and Nonnenmacher in  [DJN22]. Both [DJ18] and [DJN22] relied on the 1-dimensional fractal uncertainty principle [BD18], which restricted their results to surfaces. However, in 2023, Cohen [Co24] generalized the fractal uncertainty principle to higher dimensions. We exploit this breakthrough to examine semiclassical measures on higher-dimensional manifolds.
@@ 1.4. Structure of the paper
  • In §4, we state and generalize the higher-dimensional uncertainty principle.
@@
Acknowledgements: EK would like to thank Semyon Dyatlov for suggesting and advising this project. EK is supported by NSF GRFP under grant No. 1745302 and NM is partially supported by NSF DMS–2405264.
@@ 2.1. Hyperbolic manifolds
$X\coloneqq E_{0,1}+E_{1,0},\quad A_{k}\coloneqq E_{0,k}+E_{k,0},\quad R_{i,j}\coloneqq E_{i,j}-E_{j,i}.$
${K_{0}}$ is spanned by $R_{i+1,j+1}$ for $1\leq i,j\leq n$.
@@
We extend $\varphi_{t}$ to $T^{*}M\setminus 0$ by setting it to be homogeneous.
@@
We know that $U_{i}^{\pm}$ generate the horocyclic flows $e^{sU_{i}^{\pm}}$. In other words, if $q\in F^{*}\mathbb{H}^{n+1}$, then for all $s\in\mathbb{R}$, the geodesic starting at $\pi_{S}(e^{sU_{i}^{\pm}}q)$ has the same positive limiting point for $U_{i}^{+}$ and negative limiting point for $U_{i}^{-}$ as the geodesic starting at $\pi_{S}(q)$. For $n+1>2$, the flows of $U_{j}^{\pm}$ do not descend to $S\mathbb{H}^{n+1}$.
@@
where $E_{s}$ and $E_{u}$ are respectively the $n$-dimensional stable and unstable bundles. $E_{s}$ and $E_{u}$ were defined for $|\xi|_{g}=1$ in  [DFG15]*(3.14) by setting
@@
This definition can be extended to $|\xi|_{g}>0$ by setting $E_{s}$ and $E_{u}$ to be homogeneous. We note that $E_{s}$ and $E_{u}$ are the images of the stable and unstable bundles of $\mathbb{H}^{n+1}$ under the covering map
(2.10) $\pi_{\Gamma}:T^{*}\mathbb{H}^{n+1}\rightarrow T^{*}M.$
@@ 2.2. $U_{1}^{-}$-dense sets
A $U_{1}^{-}$-orbit is a set of the form $e^{\mathbb{R}U_{1}^{-}}q$, where $q\in F^{*}M$. A $U_{1}^{-}$-segment is a continuous subset of a $U_{1}^{-}$-orbit. Finally, we say that $U\subset S^{*}M$ is $U_{1}^{-}$-dense if $\pi^{-1}_{S}(U)$ intersects every $U_{1}^{-}$-orbit in $F^{*}M$. In other words, for each $q\in F^{*}M$, $U$ intersects $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$.
@@
Let $U$ be $U_{1}^{-}$-dense and suppose that no such $T$ exists. Then for all $j\in\mathbb{N}$, there exists $q_{j}\in F^{*}M$ such that $\pi_{S}(e^{cU_{1}^{-}}q_{j})\cap U=\emptyset$ for $c\in[-j,j]$. We pass to a convergent subsequence $q_{j}\rightarrow q_{\infty}\in F^{*}M$. Then for all $c\in\mathbb{R}$, $\pi_{S}(e^{cU_{1}^{-}}q_{\infty})\cap U=\emptyset$, a contradiction. ∎
@@
We next construct a function $F$ such that $F^{-1}(0)$ intersects a $U_{1}^{-}$-orbit in a precise way. this construction will be used for our partition of unity.
@@
    for all $q\in\pi_{S}^{-1}(\rho_{0})$, there exists $j\in\{1,\ldots,N\}$ such that $\partial_{t}^{j}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\neq 0$;
@@
Let $U\subset S^{*}M$ be open and $U_{1}^{-}$-dense. Then there exists $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$ such that $\chi_{1}+\chi_{2}=1$, $\operatorname{supp}\chi_{1}\subset U$, and the complements $S^{*}M\setminus\operatorname{supp}\chi_{1}$, $S^{*}M\setminus\operatorname{supp}\chi_{2}$ are $U_{1}^{-}$-dense.
@@
1. We first construct a particular $U_{1}^{-}$-dense set $D$. Fix $\rho_{0}\in S^{*}M$ and take $F=F_{\rho_{0}}\in C^{\infty}(U_{\rho_{0}};\mathbb{R})$ from Lemma  2.5. We know for some $j=j_{\rho_{0}}\in\{1,\ldots,N\}$, for all $q\in\pi_{S}^{-1}(\rho_{0})$, $\partial_{t}^{j}F_{\rho_{0}}(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\neq 0$. Select open $V_{\rho_{0}}^{\prime}\subset U_{\rho_{0}}$ such that for some $c>0$, for all $q\in\pi_{S}^{-1}(V_{\rho_{0}}^{\prime})$,
(2.13) $\left|\partial_{t}^{j_{\rho_{0}}}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\right|\geq c.$
Fix an open $V_{\rho_{0}}\Subset V_{\rho_{0}}^{\prime}$. Then pick $R=R_{\rho_{0}}>0$ such that for every $q\in\pi_{S}^{-1}(V_{\rho_{0}})$, $\{\pi_{S}(e^{tU_{1}^{-}}q):|t|\leq R\}\subset V_{\rho_{0}}^{\prime}$. As $F$ is smooth, for all $1\leq k\leq N$, $\partial_{t}^{k}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}$ must be uniformly bounded in $\pi^{-1}_{S}(V_{\rho_{0}})$. Additionally, $\partial_{t}^{N+1}F(\pi_{S}(e^{tU_{1}^{-}}q))$ must be uniformly bounded in $\pi^{-1}_{S}(V_{\rho_{0}})$ and $|t|\leq R$. Thus by  (2.13), we know there exists some $\varepsilon=\varepsilon_{\rho_{0}}>0$ such that for all $q\in\pi_{S}^{-1}(V_{\rho_{0}})$ and $|\tau|\leq R$
$\max_{|t|\leq R}\left|\sum_{k=1}^{N}\left(\partial_{t}^{k}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\frac{t^{k}}{k!}\right)+\partial_{t}^{N+1}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=\tau}\frac{t^{N+1}}{(N+1)!}\right|\geq\varepsilon.$
Therefore for $q\in\pi_{S}^{-1}(V_{\rho_{0}})$, by taking the Taylor expansion in $t$, we see
@@
Now, by the continuity of $F$, there exists $r>0$ such that for all $\rho\in V_{\rho_{0}}$, $|F(\rho)|<r$. Set
$D_{\rho_{0}}\coloneqq V^{\prime}_{\rho_{0}}\cap\bigcup_{\begin{subarray}{c}j\in\mathbb{Z}\\ |j|\leq\left\lceil\frac{r}{\varepsilon}\right\rceil\end{subarray}}F^{-1}(j\varepsilon).$
By (2.14), for each $q\in\pi_{S}^{-1}(V_{\rho_{0}})$, $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$ intersects $D_{\rho_{0}}$.
Now take a finite open cover of $S^{*}M$ by $V_{\rho}$, indexed by $P=\{\rho_{1},\ldots,\rho_{K}\}\subset S^{*}M$. We see that $D\coloneqq\bigcup_{\rho\in P}D_{\rho}$ is $U_{1}^{-}$-dense.
2. We now claim that $U\setminus D$ is $U_{1}^{-}$ dense. This follows from proving that for any $q\in F^{*}M$, the intersection of $D$ with $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$ has empty interior in $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$. Note that the projection onto $S^{*}M$ of each $U_{1}^{-}$-orbit has countably many segments that intersect any particular $V_{\rho}^{\prime}$. Thus, it suffices to show that for all $q\in F^{*}M$ and $t_{0}\in\mathbb{R}$, if $\pi_{S}(e^{t_{0}U_{1}^{-}}q)\in D_{\rho}$ for some $\rho\in P$, then there exists $j$ such that $\partial_{t}^{j}F_{\rho}(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=t_{0}}\neq 0$. As $e^{t_{0}U_{1}^{-}}q\in\pi_{S}^{-1}(V_{\rho}^{\prime})$, this follows from (2.13).
Therefore, for all $q\in F^{*}M$, $D\cap\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$ has empty interior in $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$. We conclude $D$ and $U\setminus D$ are both $U_{1}^{-}$-dense.
3. From Lemma  2.4, we know there exists a $U_{1}^{-}$-dense and compact $K_{1}\subset U\setminus D$. As $D\subset S^{*}M\setminus K_{1}$, we know $S^{*}M\setminus K_{1}$ is $U_{1}^{-}$-dense. Again by Lemma  2.4, we can find a $U_{1}^{-}$-dense compact $K_{2}\subset S^{*}M\setminus K_{1}$. We take a partition of unity on $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$, $\chi_{1}+\chi_{2}=1$ subordinate the the cover of $S^{*}M=(U\setminus K_{2})\cup(S^{*}M\setminus K_{1})$:

### v2

@@ Abstract.
We examine semiclassical measures for Laplace eigenfunctions on compact hyperbolic $(n+1)$-manifolds. We prove their support must contain the cosphere bundle of a compact immersed totally geodesic submanifold of dimension at least 2. Our proof adapts the argument of [DJ18] and [ADM24] and classifies the closures of unipotent orbits using Ratner theory. An important step in the proof is a generalization of the higher-dimensional fractal uncertainty principle of Cohen [Co24] to Fourier integral operators, which may be of independent interest.
@@ 1.1. Main Results
Let $U_{1}^{-}$ be the vector field on $F^{*}M$ defined in (2.6). This vector field is the generator of the one-parameter unipotent flow $e^{sU_{1}^{-}}$.
It is well-known (see for example  [Zw12]*§§5.1–2) that if $\mu$ is a semiclassical measure, $\mu$ is invariant under the geodesic flow $\varphi_{t}:S^{*}M\rightarrow S^{*}M$ and $\operatorname{supp}\mu\subset S^{*}M$.
@@
As discussed in §3.6, there are many 3-dimensional hyperbolic manifolds with no totally geodesic surfaces. On any such manifold, Theorem 1.2 implies semiclassical measures must have full support.
We remark here that the presence of cosphere bundles of totally geodesic submanifolds in the higher-dimensional statement of both Theorem 1.2 and the main theorem of [ADM24] is intimately related to the failure of one-parameter unipotent flows on $F^{*}M$ to be uniquely ergodic. This should be juxtaposed with the $2$-dimensional case [DJ18], where the unique ergodicity of the horocycle flow is a key tool in concluding that semiclassical measures have full support.
@@ 1.2. Previous results
Developed in the 1970s and 80s, the Quantum Ergodicity theorem of Shnirelman [Sh74], Zelditch [Ze87], and Colin de Verdière [CdV85] implies that on negatively curved compact manifolds, a density-one sequence of Laplace eigenfunctions converges semiclassically to the Liouville measure.
@@
Theorems 1.1 and 1.2 exclude different semiclassical measures than the entropy bounds. This distinction is perhaps easiest to state in dimension $3$, where the existence of infinitely many commensurability classes of compact hyperbolic $3$-manifolds with no immersed, totally geodesic surfaces is known [MRBook, Theorem 9.5.1]. For such examples, Theorem 1.2 asserts that any semiclassical measure has full support, in contrast to the abundance of measures with entropies greater than $1$ with proper support. In higher dimensions, one can make similar statements, though a limiting factor is our understanding of the behavior of geodesic submanifolds in non-arithmetic manifolds (see §3.6). However, the entropy bounds exclude some measures that our results do not. One example is a semiclassical measure of the form $\mu=\alpha\mu_{L}+(1-\alpha)\mu_{0}$, where $\mu_{L}$ is the Liouville measure, $\mu_{0}$ is a $\delta$-measure on a closed geodesic, and $20<\alpha<1/$ 2. Such a measure clearly has full support, but has entropy $\alpha n<n/2$, and therefore is ruled out by [AN07].
Our work, specifically Theorem 1.1, builds off of [DJ18], wherein Dyatlov and Jin proved that all semiclassical measures on compact hyperbolic surfaces $M=\Gamma\backslash\mathbb{H}^{2}$ have full support. This result was generalized to compact surfaces with negative curvature by Dyatlov, Jin, and Nonnenmacher in  [DJN22]. Both [DJ18] and [DJN22] relied on the 1-dimensional fractal uncertainty principle [BD18], which restricted their results to surfaces. However, in 2023, Cohen [Co24] generalized the fractal uncertainty principle to higher dimensions. His proof used the work of Han and Schlag from [HS20]. We exploit this breakthrough to examine semiclassical measures on higher-dimensional manifolds.
@@ 1.4. Structure of the paper
  • In §4, we state and generalize the higher-dimensional fractal uncertainty principle.
@@
Acknowledgements: EK would like to thank Semyon Dyatlov (partially supported by NSF–DMS-2400090 and NSF CAREER grant DMS–1749858) for suggesting and advising this project. EK is supported by NSF GRFP under grant No. 1745302 and NM is partially supported by NSF DMS–2405264.
@@ 2.1. Hyperbolic manifolds
$X\coloneqq E_{0,1}+E_{1,0},\quad A_{k}\coloneqq E_{0,k}+E_{k,0},\quad R_{i,j}\coloneqq E_{i,j}-E_{j,i},$
where $E_{i,j}$ is the elementary matrix with all entries zero, except for the $(i,j)$th entry, which is equal to 1. ${K_{0}}$ is spanned by $R_{i+1,j+1}$ for $1\leq i,j\leq n$.
@@
We extend $\varphi_{t}$ to $T^{*}M\setminus 0$ by setting it to be homogeneous. By homogeneity, we mean $[X,\xi\cdot\partial_{\xi}]=0$, where $\xi\cdot\partial_{\xi}$ is the generator of dilations.
@@
We know that $U_{i}^{\pm}$ generate the unipotent flows $e^{sU_{i}^{\pm}}$. In other words, if $q\in F^{*}\mathbb{H}^{n+1}$, then for all $s\in\mathbb{R}$, the geodesic starting at $\pi_{S}(e^{sU_{i}^{\pm}}q)$ has the same positive limiting point for $U_{i}^{+}$ and negative limiting point for $U_{i}^{-}$ as the geodesic starting at $\pi_{S}(q)$.
@@
where $E_{s}$ and $E_{u}$ are respectively the $n$-dimensional stable and unstable bundles. These stable and unstable bundles were defined over $S^{*}\mathbb{H}^{n+1}$ in  [DFG15]*(3.14) by setting
@@
This definition can be extended to $|\xi|_{g}>0$ by setting $E_{s}$ and $E_{u}$ to be homogeneous.
@@
In $T_{(x,\xi)}(T^{*}M)$, $E_{s}$ and $E_{u}$ are the images of the stable and unstable bundles of $\mathbb{H}^{n+1}$ under the covering map
(2.11) $\pi_{\Gamma}:T^{*}\mathbb{H}^{n+1}\rightarrow T^{*}M.$
@@ 2.2. $U_{1}^{-}$-dense sets
  • A $U_{1}^{-}$-orbit is a set of the form $e^{\mathbb{R}U_{1}^{-}}q$, where $q\in F^{*}M$.
  • A $U_{1}^{-}$-segment is a set of the form $\{e^{tU_{1}^{-}}q:a\leq t\leq b\}$, where $q\in F^{*}M$ and $a<b\in\mathbb{R}$.
  • A set $U\subset S^{*}M$ is $U_{1}^{-}$-dense if $\pi^{-1}_{S}(U)$ intersects every $U_{1}^{-}$-orbit in $F^{*}M$. In other words, for each $q\in F^{*}M$, $U$ intersects $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$.
  • Finally, a set $U\subset S^{*}M$ is $U_{1}^{-}$-sparse if for all $q\in F^{*}M$, $\{t\in\mathbb{R}:\pi_{S}(e^{tU_{1}^{-}}(q))\in U\}$ is countable.
Clearly, if $U$ is $U_{1}^{-}$-sparse, then $S^{*}M\setminus U$ is $U_{1}^{-}$-dense.
@@
Let $U$ be $U_{1}^{-}$-dense and suppose that no such $T$ exists. Then for all $j\in\mathbb{N}$, there exists $q_{j}\in F^{*}M$ such that $\pi_{S}(e^{cU_{1}^{-}}q_{j})\cap U=\emptyset$ for all $c\in[-j,j]$. We pass to a convergent subsequence $q_{j}\rightarrow q_{\infty}\in F^{*}M$. Then for all $c\in\mathbb{R}$, $\pi_{S}(e^{cU_{1}^{-}}q_{\infty})\cap U=\emptyset$, a contradiction. ∎
@@
We next construct a function $F$ which has a particular relationship with $U_{1}^{-}$-orbits. This relationship will be used to build a set $D\subset S^{*}M$ that is both $U_{1}^{-}$-sparse and $U_{1}^{-}$-dense, which in turn will be used to create a partition of unity in Lemma 2.7. Lemma 2.7 is a key step in the proof of Lemma 5.1.
@@
    for each $q\in\pi_{S}^{-1}(\rho_{0})$, there exists $j=j_{q}\in\{1,\ldots,N\}$ such that $\partial_{t}^{j}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\neq 0$;
@@
Lemma 2.6.
Let $\rho_{0}\in S^{*}M$ and let $F\in C^{\infty}(U_{\rho_{0}})$ be the function constructed in Lemma 2.5. There exists some $U^{\prime}_{\rho_{0}}\subset U_{\rho_{0}}$ and $c=c_{\rho_{0}}>0$ such that for each $q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime})$, for some $j_{q}\in\{1,\ldots,N\}$,
(2.13) $\left|\partial_{t}^{j_{q}}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\right|\geq c.$
Additionally, for any $r\in\mathbb{R}$, $F^{-1}(r)\cap U^{\prime}_{\rho_{0}}$ is $U_{1}^{-}$-sparse.
Proof.
For each $q\in\pi_{S}^{-1}(\rho_{0})$, there exists $j_{q}\in\{1,\ldots,N\}$ such that $\partial_{t}^{j_{q}}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\neq 0$. Thus, for each $q\in\pi_{S}^{-1}(\rho_{0})$, we can select open $U_{q}\subset\pi_{S}^{-1}(U_{\rho_{0}})$ such that $q\in U_{q}$ and for some $c_{q}>0$, if $\tilde{q}\in U_{q}$, then $|\partial_{t}^{j_{q}}F(\pi_{S}(e^{tU_{1}^{-}}\tilde{q}))|_{t=0}|\geq c_{q}$. By the compactness of $\pi_{S}^{-1}(\rho_{0})$, we know
$U^{\prime}_{\rho_{0}}\coloneqq\bigcap_{q\in\pi^{-1}_{S}(\rho_{0})}\pi_{S}(U_{q})\subset U_{\rho_{0}}$
is a nonempty open set and $c\coloneqq\inf_{q\in\pi_{S}^{-1}(\rho_{0})}c_{q}>0$. Therefore, for each $q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime})$, there exists some $j_{q}\in\{1,\ldots,N\}$ such that
$\left|\partial_{t}^{j_{q}}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\right|\geq c,$
which completes the proof of  (2.13).
Now we show for any $r\in\mathbb{R}$, $F^{-1}(r)\cap U_{\rho_{0}}^{\prime}$ is $U_{1}^{-}$-sparse. Note that the projection onto $S^{*}M$ of each $U_{1}^{-}$-orbit has countably many segments that intersect $F^{-1}(r)\cap U_{\rho_{0}}^{\prime}$. Thus, it suffices to show that for all $q\in F^{*}M$ and $t_{0}\in\mathbb{R}$, if $\pi_{S}(e^{t_{0}U_{1}^{-}}q)\in F^{-1}(r)\cap U_{\rho_{0}}^{\prime}$, then there exists $j$ such that $\partial_{t}^{j}F_{\rho}(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=t_{0}}\neq 0$. As $e^{t_{0}U_{1}^{-}}q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime})$, this follows from (2.13). ∎
@@
Let $U\subset S^{*}M$ be open and $U_{1}^{-}$-dense. Then there exist $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$ such that $\chi_{1}+\chi_{2}=1$, $\operatorname{supp}\chi_{1}\subset U$, and the complements $S^{*}M\setminus\operatorname{supp}\chi_{1}$, $S^{*}M\setminus\operatorname{supp}\chi_{2}$ are $U_{1}^{-}$-dense.
@@
1. We first use Lemma 2.6 to construct a set $D\subset S^{*}M$ that is both $U_{1}^{-}$-dense and $U_{1}^{-}$-sparse. Fix $\rho_{0}\in S^{*}M$ and take $F=F_{\rho_{0}}\in C^{\infty}(U_{\rho_{0}};\mathbb{R})$ from Lemma  2.5 and $U_{\rho_{0}}^{\prime}\subset U_{\rho_{0}}$ and $c_{\rho_{0}}$ from Lemma  2.6. Fix an open $U_{\rho_{0}}^{\prime\prime}\Subset U_{\rho_{0}}^{\prime}$. Then pick $R=R_{\rho_{0}}>0$ such that for every $q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime\prime})$, $\{\pi_{S}(e^{tU_{1}^{-}}q):|t|\leq R\}\subset U_{\rho_{0}}^{\prime}$. We assume that $R\leq 1$.
Let $q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime\prime})$. By (2.13), we can define
$P_{q}(t)=\left(\partial_{t}^{j_{q}}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\right)^{-1}\sum_{k=1}^{j_{q}}\partial_{t}^{k}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\frac{t^{k}}{k!}.$
Using Chebyshev polynomials, one can show that among monic polynomials of degree $n$, the smallest maximal absolute value on the interval $[-1,1]$ is $2^{1-n}$. Therefore, we know that $\max_{|t|\leq R}|P_{q}(t)|\geq 2^{1-j_{q}}\geq 2^{1-N}$.
As $F$ is smooth, for $|t|\leq R$ and $11\leq k\leq N+$ 1, $|\partial_{t}^{k}F(\pi_{S}(e^{tU_{1}^{-}}q))|\leq C$, where $C$ is uniform in $t$, $k$, and $q\in\pi^{-1}_{S}(U_{\rho_{0}}^{\prime\prime})$. Therefore, taking the Taylor expansion in $t$ and using (2.13),
$\max_{|t|\leq R}|F(\pi_{S}(e^{tU_{1}^{-}}q))-F(\pi_{S}(q))|$ $\geq\max_{|t|\leq R}\left(\left|\sum_{k=1}^{j_{q}}\partial_{t}^{k}F(\pi_{S}(e^{tU_{1}^{-}}q))|_{t=0}\frac{t^{k}}{k!}\right|-C\frac{|t|^{j_{q}+1}}{(j_{q}+1)!}\right)$
$\geq c_{\rho_{0}}2^{1-N}-CR.$
Possibly shrinking the value of $R$, there exists $\varepsilon=\varepsilon_{\rho_{0}}>0$ such that $c_{\rho_{0}}2^{1-N}-CR\geq\varepsilon$. We conclude
@@
Now, by the continuity of $F$, there exists $r>0$ such that for all $\rho\in U_{\rho_{0}}^{\prime\prime}$, $|F(\rho)|<r$. Set
$D_{\rho_{0}}\coloneqq U^{\prime}_{\rho_{0}}\cap\bigcup_{\begin{subarray}{c}j\in\mathbb{Z}\\ |j|\leq\left\lceil\frac{r}{\varepsilon}\right\rceil\end{subarray}}F^{-1}(j\varepsilon).$
By (2.14), for each $q\in\pi_{S}^{-1}(U_{\rho_{0}}^{\prime\prime})$, $\pi_{S}(e^{\mathbb{R}U_{1}^{-}}q)$ intersects $D_{\rho_{0}}$.
Now take a finite open cover of $S^{*}M$ by $U_{\rho}^{\prime\prime}$, indexed by $P=\{\rho_{1},\ldots,\rho_{K}\}\subset S^{*}M$. We see that $D\coloneqq\bigcup_{\rho\in P}D_{\rho}$ is $U_{1}^{-}$-dense.
By Lemma 2.6, $D$ is the finite union of $U_{1}^{-}$-sparse sets. Therefore, $D$ is $U_{1}^{-}$-sparse.
2. Note that both $D$ and $U\setminus D$ are $U_{1}^{-}$-dense. From Lemma  2.4, we know there exists a $U_{1}^{-}$-dense and compact $K_{1}\subset U\setminus D$. As $D\subset S^{*}M\setminus K_{1}$, we know $S^{*}M\setminus K_{1}$ is $U_{1}^{-}$-dense. Again by Lemma  2.4, we can find a $U_{1}^{-}$-dense compact $K_{2}\subset S^{*}M\setminus K_{1}$. We take a partition of unity on $\chi_{1},\chi_{2}\in C^{\infty}(S^{*}M;[0,1])$, $\chi_{1}+\chi_{2}=1$ subordinate the the cover of $S^{*}M=(U\setminus K_{2})\cup(S^{*}M\setminus K_{1})$: